Active Adventures will pay an annual dividend of RM3.15 a share on their common stock
next week. Last year, the company paid a dividend of RM3.00 a share. The company
adheres to a constant rate of growth dividend policy. What will one share of this common
stock be worth ten years from now if the applicable discount rate is 12.5 percent?
1
Expert's answer
2016-05-06T12:16:03-0400
Annual dividend of RM3.15 a share on their common stock, last year dividend of RM3.00 a share. The company adheres to a constant rate of growth dividend policy. What will one share of this common stock be worth ten years from now if the applicable discount rate is 12.5 percent?
A constant growth stock is a stock whose dividends are expected to grow at a constant rate in the forseeable future. This condition fits many established firms, which tend to grow over the long run at the same rate as the economy, fairly well. The value of a constant growth stock can be determined using the following equation:
P0 = D0*(1 + g)/(r - g) = D1/(r - g), where P0 = the stock price at time 0, D0 = the current dividend, D1 = the next dividend (i.e., at time 1), g = the growth rate in dividends, and r = the required return on the stock, and g < r.
g = D1/D0 - 1 = 3.15/3 - 1 = 0.05.
So, P0 = 3.15*(1 + 0.05)/(0.125 - 0.05) = $44.1 per share.